# Planetarium[Uncaptioned image]: A Rigorous Benchmark for Translating Text to Structured Planning Languages

### v1

@@ Abstract
Many recent works have explored using language models for planning problems. One line of research focuses on translating natural language descriptions of planning tasks into structured planning languages, such as the planning domain definition language (PDDL). While this approach is promising, accurately measuring the quality of generated PDDL code continues to pose significant challenges. First, generated PDDL code is typically evaluated using planning validators that check whether the problem can be solved with a planner. This method is insufficient because a language model might generate valid PDDL code that does not align with the natural language description of the task. Second, existing evaluation sets often have natural language descriptions of the planning task that closely resemble the ground truth PDDL, reducing the challenge of the task. To bridge this gap, we introduce Planetarium, a benchmark designed to evaluate language models’ ability to generate PDDL code from natural language descriptions of planning tasks. We begin by creating a PDDL equivalence algorithm that rigorously evaluates the correctness of PDDL code generated by language models by flexibly comparing it against a ground truth PDDL. Then, we present a dataset of $132,037$ text-to-PDDL pairs across $13$ different tasks, with varying levels of difficulty. Finally, we evaluate several API-access and open-weight language models that reveal this task’s complexity. For example, $87.6\%$ of the PDDL problem descriptions generated by GPT-4o are syntactically parseable, $82.2\%$ are valid, solve-able problems, but only $35.1\%$ are semantically correct, highlighting the need for a more rigorous benchmark for this problem.
@@ 1 Introduction
Recently, there has been growing interest in using large language models (LLMs) to solve planning problems. Some research has focused on generating plans directly with LLMs Valmeekam et al. (2023a, 2022, b); Silver et al. (2022, 2023). However, this approach has shown limited success; GPT-4 only achieves $35\%$ accuracy on simple planning problems Valmeekam et al. (2023a). Another line of research uses LLMs to convert natural language prompts into structured planning languages, such as the planning domain definition language (PDDL) Liu et al. (2023); Xie et al. (2023b); Guan et al. (2023); Chalvatzaki et al. (2023); Yang et al. (2023b). Early evidence suggests this method performs better than generating plans directly with LLMs Liu et al. (2023). Despite its promise, there is a lack of rigorous techniques and benchmarks for evaluating the translation of natural language planning descriptions to PDDL.
@@
Using LLMs to translate natural language to PDDL is an instance of a code generation task. Evaluating code generation tasks in general is highly challenging. Some benchmarks for code generation use match-based metrics that look for segments that overlap with ground truth code Papineni et al. (2002); Lin (2004); Ren et al. (2020), but match-based metrics cannot cover the vast space of equivalent programs Chen et al. (2021). Therefore, benchmarks often test functional correctness using a suite of unit tests Roziere et al. (2020); Kulal et al. (2019). For planning problems, existing works use “plan validators” to check if the generated code can be solved with a traditional planner Liu et al. (2023); Silver et al. (2022, 2023); Guan et al. (2023). We argue that these validators alone are insufficient to determine if PDDL generation is correct. This is because the LLM can generate valid PDDL that has nothing to do with the user’s instructions and still be considered correct.
Figure 1: An illustration of how multiple PDDL problem descriptions can describe the same planning problem. All four PDDL problem descriptions represent the planning problem of stacking one block onto another.
A rigorous evaluation of LLMs as generators of structured planning languages requires a precise definition of what it means for generated code to be correct, and an efficient, automatic way of checking it. This requirement is hard because many instances of PDDL can represent the same planning problem, but it is not always obvious when they are equivalent (Figure 1). Properly checking PDDL equivalence requires symbolic interpretation.
@@
  • Rigorous Evaluation of PDDL Equivalence. We formally define planning problem equivalence and create an algorithm for checking whether two PDDL problems satisfy this definition. This algorithm transforms the PDDL code into scene graphs, computes an expansion of the goal states for both PDDL problems, and then performs isomorphism checks between the graphs. Our method ensures two PDDL problems match if and only if they represent the same underlying planning problem. Further, we show how to make this algorithm practically efficient for two domains: Blocks World and Gripper McDermott (2000) (Appendix C).
  • Benchmark Data for PDDL Generation. We present a dataset based on the International Planning Competition (IPC) McDermott (2000); Vallati et al. (2015); Taitler et al. (2024); Seipp et al. (2022), crafting $132,037$ ground truth PDDL problems and corresponding text descriptions capturing a range of planning problems. Each task varies in two dimensions to assess the difficulty of the PDDL generation task: abstraction and size.
  • Broad Evaluation of Current LLMs. Finally, we evaluate a range of API-access and open-access LLMs on Planetarium. We evaluate in both a zero-shot setting and after fine-tuning. We find that this task is very challenging. GPT-4o in a zero-shot setting gets only $35.1\%$ correct. Instances with abstract descriptions or many propositions are particularly challenging. Planetarium can therefore serve as a benchmark of progress on this important problem. To enable future development and evaluation of LLMs we release all the code and data for Planetarium at https://github.com/BatsResearch/planetarium.
@@ 2 Related Work
LLMs as Planners.
There is growing interest in using LLMs for solving planning problems. Several works have demonstrated the potential of using LLMs for decision making processes Sharma et al. (2022); Ren et al. (2023); Ichter et al. (2022); Singh et al. (2022). Some techniques that enable the use of LLMs as planners involve decomposing abstract tasks into concrete, atomic executable actions for an agent to perform Huang et al. (2022b, a); Sharma et al. (2022). Other approaches generate plan actions by scoring the possible next steps in a sequence of actions Ichter et al. (2022); Ren et al. (2023). When using LLMs to generate plans directly, the LLM is given a natural language representation of a planning problem with the goal of generating a plan. We refer to this line of work as “LLMs as planners.” One of the main findings is that LLMs have limited ability to generate and validate plans on their own, even for simple planning tasks Valmeekam et al. (2022); Silver et al. (2022); Valmeekam et al. (2023b).
Planner-Augmented LLMs. Given LLMs’ poor performance in classical planning tasks Valmeekam et al. (2023a, 2022); Silver et al. (2022), new approaches extend LLMs by incorporating classical planning techniques Liu et al. (2023); Guan et al. (2023). Some techniques frame the problem as a machine translation task, transforming a natural language text description of a planning problem into PDDL to use with classical planners Liu et al. (2023); Guan et al. (2023); Xie et al. (2023a). We refer to this line of work as “Planner-Augmented LLMs.” Other similar approaches translate natural language into alternative representations of the planning problem, such as finite state automata Yang et al. (2023a) or logic programming Yang et al. (2023b), to solve them.
Benchmarking LLMs on Planning Tasks. Significant efforts have been made to develop benchmarks to assess the use of “LLMs as Planners.” PlanBench Valmeekam et al. (2023a) evaluates LLMs on various planning-related tasks, including plan generation, cost-optimal planning, plan verification, and other properties. Their work focuses on evaluating LLMs on their ability to generate plans. In contrast Planetarium focuses on evaluating “Planner-Augmented LLMs”. Here, the goal is to determine if an LLM can successfully translate natural language descriptions of planning problems into the correct PDDL representations of those problems.
@@ 3.1 Classical Planning Problems
In a classical planning problem, states are fully observable, there is a finite set of actions, and transitions are deterministic. The goal is to transition from the initial state to the goal state following a transition function. We use the set-theoretic definition of classical planning problems Ghallab et al. (2004).
@@
  • $A$ represents the set of actions, where each action $a$ consists of a set of preconditions that must hold for it to be applicable and a set of effects that modify the propositions that are true after its execution.
  • The transition function $\gamma:S\times A\to 2^{S}$ models how the world changes by an action.
@@
A plan $\pi=<a_{1},a_{n},\ldots,a_{n}>$ is a sequence of actions that leads from $s_{i}$ to any state in $S_{g}$ following $\gamma$. This plan is the solution to a planning problem $P$.
@@ 3.2 Planning Domain Definition Language
The Planning Domain Definition Language (PDDL) is a specialized language designed to provide a unified way to represent planning problems. It can represent various types of planning problems, including classical planning problems McDermott (2000); Gerevini and Long (2006). A PDDL planning problem consists of two files: the domain file, which describes the constant parts of the world model, and a problem file, which specifies particular planning scenarios using the world model outlined in the domain file.
In the domain file, we specify the set of possible actions $A$ and their preconditions and effects, which collectively define the transition function $\gamma$. The problem file defines the initial state $s_{i}$ and the goal propositions $g$. Finally, the set of proposition symbols $L$ is created by combining the predicates from the domain file with the objects defined in the problem file.
@@ 3.3 Scene Graphs
To compare different planning problems, we use scene graphs. A scene graph is a data structure commonly used in fields such as computer vision and graphics Johnson et al. (2015); Chang et al. (2023), rearrangement Ramachandruni et al. (2023), and planning to represent objects, their attributes, and the relationships among them. In our work, we define scene graphs as directed graphs with types and attributes for both nodes and edges. We represent a PDDL problem file with scene graphs as follows. We create one scene graph (the initial scene) for the initial state and another for the set of goal propositions (the goal scene). For every object, we create a node with an object type. Then for every proposition that is listed in the problem file, we create a node with a proposition type. That node is given an attribute with the name of its predicate. Then, for each argument to the predicate in that proposition, we add an edge from the proposition node to the corresponding object. Each such edge is given three attributes: the name of the predicate, the position of the argument (first, second, etc.), and whether it is defined in the initial state or the goal propositions. A scene graph is thus ($(\{O\cup P\},E)$, where $O$ is the set of nodes with object type, $P$ is the set of nodes with Proposition type, and $E$ is the set of edges.
We further define a problem graph as the combination of an initial scene and goal scene. Given $SceneGraph_{init}=(\{O\cup P_{init}\},E_{init})$ and $SceneGraph_{goal}=(\{O\cup P_{goal}\},E_{goal})$, then a problem graph merges the scene graphs such that $ProblemGraph=(\{O\cup P_{init}\cup P_{goal}\},\{E_{init}\cup E_{goal}\})$. We define graph isomorphism on any of these graphs as an edge and type-preserving bijection between nodes, meaning that two graphs share a connectivity structure where all types and attributes match. See Appendix B for diagrams of scene and problem graphs.
@@ 4 Evaluation Framework
In this section, we describe the design of Planetarium. Our goal is to systematically evaluate how well LLMs can translate natural language descriptions of planning tasks into structured planning languages. Planetarium consists of two components: an algorithm that validates whether a ground truth PDDL problem file and an LLM-generated PDDL problem file are equivalent, and a curated dataset of planning tasks against which an LLM can be evaluated.
@@ 4.1 Planning Problem Equivalence
The first step to benchmarking PDDL generation is determining how to decide whether the generated code matches ground truth code. One might assume that checking if two PDDL problem files are equivalent is straightforward. However, the same initial state, goal state pair could be represented by many PDDL problem files, as shown in Figure 1.
Given these difficulties, we propose a definition of equivalence in terms of classical planning problems. The main idea is to find a bijective function between the sets of proposition symbols $L$ of each problem, such that when applied to the other elements, this function makes both problems equal. This definition assumes that the set of actions $A$ and the transition function $\gamma$ are shared between the two problems. Our formal definition of equivalence between two planning problems is:
@@
Two planning problems $P_{1}=(L^{1},S^{1},A,\gamma,s_{i}^{1},g^{1})$ and $P_{2}=(L^{2},S^{2},A,\gamma,s_{i}^{2},g^{2})$ with the same actions $A$ and transition function $\gamma$ are equivalent if there exists a bijective function $f:L^{1}\to L^{2}$ such that:
@@
This definition is not directly usable for checking equivalence between two PDDL problem files. This is because our definition relies on finding a bijection between the sets $L$ of each problem, and PDDL does not define these sets directly. To build the set $L$ with PDDL, one needs to use the predicates in the domain file and the objects in the problem file, instantiating each predicate with all possible objects they might take. Since the set of actions $A$ and the transition function $\gamma$ are defined in the PDDL domain file, and since the equivalence definition assumes they are shared, we can assume the entire PDDL domain is shared. This fact entails that the predicates will be shared, making it necessary to look for bijective functions for PDDL only over the objects. The challenge is that each PDDL problem file can correspond to many pairs of initial and goal states because the set of goal propositions makes an open-world assumption. That set can leave implicit trivial propositions that are necessarily true and therefore do not change the underlying planning problem. We must therefore fully specify the goal, meaning that we identify all propositions that are true in all reachable goal states when starting from the initial state.
Our algorithm for checking equivalence is summarized in Algorithm 1. First, we transform each set of initial state and goal state propositions into scene graphs. Second, we fully specify the goal scene graphs by adding all trivially true edges. Finally, we join the initial state scene graph with each goal state graph to create problem graphs and look for a bijection between objects such that the problem graphs are isomorphic. We now discuss these components in more detail.
@@
1:function isEquivalent($\text{P}_{a},\text{P}_{b},\text{isPlaceholder}$)
@@
4:
@@
6:         return $\text{fastEquivalence}(\text{s}_{i,b},\text{g}_{a},\text{s}_{i,a},\text{g}_{b})$
7:     end if
8:
9:     $\text{g}^{\star}_{a}\leftarrow\text{fullySpecify}(\text{s}_{i,a},\text{g}_{a})$
10:     $\text{g}^{\star}_{b}\leftarrow\text{fullySpecify}(\text{s}_{i,b},\text{g}_{b})$
11:
12:     if isPlaceholder then
13:         return $\text{isIsomorphic}(\text{s}_{i,a},\text{s}_{i,b})\wedge\text{isIsomorphic}(\text{g}^{\star}_{a},\text{g}^{\star}_{b})$
14:     else
15:         $\text{p}_{a}\leftarrow\text{join(}\text{s}_{i,a},\text{g}^{\star}_{a}\text{)}$
16:         $\text{p}_{b}\leftarrow\text{join(}\text{s}_{i,b},\text{g}^{\star}_{b}\text{)}$
17:
18:         return $\text{isIsomorphic}(\text{p}_{a},\text{p}_{b})$
19:     end if
20:end function
@@
Transform to Scene Graphs. From each PDDL problem file, we generate two scene graphs: one for the initial state and another for the goal state (lines 2–3). For each transformation, we first create an object node for each object. Then, for each proposition in the collection, we create a new proposition node and create edges between the proposition node and the objects it takes as arguments. Edge attributes denote argument order, predicate type, and whether the proposition is in an initial or goal scene. See Section 3.3 for further details.
@@
Check Easy Cases. For speed, we check several cases and return early if we can (lines 5–7). First, if the number of object nodes in each graph is not equal, the problems cannot be equivalent. Second, if the initial scenes are not isomorphic, then the problems cannot be equivalent. Finally, if the problem graphs, composed of the initial and goal scenes, are isomorphic, then the problems are equivalent.
@@
Fully Specify the Goal Scenes. If the input is not an easy case, then we have to reason about the sets of goal states defined by the goal scenes. Condition 3 of Definition 2 requires that the sets of goal states that are consistent with the given goal propositions be equal after substituting matching propositions. Since PDDL uses the open-world assumption for goals, we have to identify all propositions that are true in all reachable goal states when starting from the initial state for each problem. The function fullySpecify finds all such propositions and adds them to the goal scenes as additional proposition nodes with the corresponding edges and attributes (lines 9–10). We show how to implement fullySpecify efficiently for Blocks World and Gripper in Appendix C.
@@
PDDL without Object Identity. Our algorithm runs in two modes, depending on the type of problem we are checking. Sometimes we want to compare a generated PDDL problem file with many ground truth problem files and see if the generated file matches any one of them. For example, if the natural language description says to “make a tower of height 3,” the specific blocks to use are unspecified, and any permutation of blocks that builds a tower of height 3 should be considered correct. Concretely, we want to treat the objects in the PDDL goal states as placeholders and accept any permutation of them. We check this condition when isPlaceholder is True (determined by the problem type) simply by checking isomorphism between initial and goal scenes separately rather than combining them into problem graphs (line 13).
@@
PDDL with Object Identity. If isPlaceholder is False then we want to check whether the problem files are precisely equivalent under Defintion 2, meaning that the objects in the initial scenes correspond to the same objects in the goal scenes as well. We first join the corresponding initial and scene graphs into problem graphs as described in Section 3.3 (lines 15–16). Then we check whether the two problem graphs are isomorphic (line 18).
@@
isEquivalent($P_{a}$, $P_{b}$, False) returns True if and only if the PDDL problem files $P_{a}$ and $P_{b}$ represent equivalent planning problems under Definition 2. isEquivalent($P_{a}$, $P_{b}$, True) returns True if and only if $P_{a}$ represents a planning problem that is equivalent to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state.

### v2

@@ Abstract
Recent works have explored using language models for planning problems. One approach examines translating natural language descriptions of planning tasks into structured planning languages, such as the planning domain definition language (PDDL). Existing evaluation methods struggle to ensure semantic correctness and rely on simple or unrealistic datasets. To bridge this gap, we introduce Planetarium, a benchmark designed to evaluate language models’ ability to generate PDDL code from natural language descriptions of planning tasks. Planetarium features a novel PDDL equivalence algorithm that flexibly evaluates the correctness of generated PDDL, along with a dataset of 145,918 text-to-PDDL pairs across 73 unique state combinations with varying levels of difficulty. Finally, we evaluate several API-access and open-weight language models that reveal this task’s complexity. For example, 96.1% of the PDDL problem descriptions generated by GPT-4o are syntactically parseable, 94.4% are solvable, but only 24.8% are semantically correct, highlighting the need for a more rigorous benchmark for this problem.
Planetarium[Uncaptioned image]: A Rigorous Benchmark for
Translating Text to Structured Planning Languages
Max Zuothanks: These authors contributed equally to this work.  Francisco Piedrahita Velez 11footnotemark: 1  Xiaochen Li Michael L. LittmanStephen H. Bach 1Department of Computer Science, Brown University {zuo, francisco, xiaochen_li, michael_littman, stephen_bach}@brown.edu
@@ 1 Introduction
Figure 1: An example of one planning goal corresponding to many correct PDDL goals. All PDDL goals in the top row represent the displayed goal correctly. The bottom row illustrates PDDL goals with different error types, showing instances that are solvable (a planner can generate a plan, but for a different planning problem), parseable (the PDDL syntax is correct but will not produce any plan from a planner), and not parseable (it is not valid PDDL). See Section 5 for details.
Recently, there has been growing interest in using large language models (LLMs) to solve planning problems. Some research has focused on generating plans directly with LLMs Valmeekam et al. (2023a, 2022, b); Silver et al. (2022, 2023). However, this approach has shown limited success; GPT-4 only achieves 35% accuracy on simple planning problems Valmeekam et al. (2023a). Another line of research uses LLMs to convert natural language prompts into structured planning languages, such as the planning domain definition language Liu et al. (2023); Xie et al. (2023); Guan et al. (2023); Chalvatzaki et al. (2023); Yang et al. (2023b). Early evidence suggests this method performs better than generating plans directly with LLMs Liu et al. (2023). Despite its promise, there are no rigorous techniques or benchmarks for evaluating the translation of natural language planning descriptions to PDDL.
@@
Using LLMs to translate natural language to PDDL is an instance of a code generation task. Evaluating code generation tasks, in general, is highly challenging. Some benchmarks for code generation use match-based metrics that look for segments that overlap with ground truth code Papineni et al. (2002); Lin (2004); Ren et al. (2020), but match-based metrics cannot cover the vast space of equivalent programs Chen et al. (2021). Therefore, benchmarks often test functional correctness using a suite of unit tests Roziere et al. (2020); Kulal et al. (2019). For planning problems, existing work uses “plan validators” to check if the generated code can be solved with a traditional planner Liu et al. (2023); Silver et al. (2022, 2023); Guan et al. (2023). We argue using validators alone is insufficient to determine if PDDL generation is correct. This is because the LLM can generate valid PDDL that has nothing to do with the user’s instructions and still be considered correct—a false positive.
A rigorous evaluation of LLMs as generators of structured planning languages requires a precise definition of what it means for generated code to be correct. This is hard because many instances of PDDL can represent the same planning problem, but it is not always obvious when they are equivalent (Figure 1). Properly checking PDDL equivalence requires symbolic interpretation.
@@
Rigorous Evaluation of PDDL Equivalence.
We formally define planning problem equivalence and create an algorithm for checking whether two PDDL problems satisfy this definition. This algorithm transforms the PDDL code into scene graphs, computes an expansion of the goal states for both PDDL problems, and then performs isomorphism checks between the graphs. Our method ensures two PDDL problems match if and only if they represent the same underlying planning task. We show how to make this algorithm efficient for three domains: Blocks World, Gripper, and a slightly simplified Floor Tile McDermott (2000); Linares López et al. (2015).
Benchmark Data for PDDL Generation.
We present a dataset based on the International Planning Competition (IPC) McDermott (2000); Vallati et al. (2015); Taitler et al. (2024); Seipp et al. (2022), crafting 145,918 ground truth PDDL problems and corresponding text descriptions capturing a range of planning problems. Each task varies in two dimensions to assess the difficulty of PDDL generation: abstraction and size.
Broad Evaluation of Current LLMs.
Finally, we evaluate a range of API-access and open-weight LLMs on Planetarium. We evaluate in both a zero-shot setting and after fine-tuning. We find that this task is very challenging. GPT-4o in a zero-shot setting gets only 24.8% correct. Instances with abstract descriptions or many propositions are particularly challenging. Planetarium can, therefore, serve as a benchmark of progress on this important problem. To support future development and evaluation of LLMs, we release all the code and data for Planetarium at github.com/BatsResearch/planetarium.
@@ 2 Related Work
LLMs as Planners. There is growing interest in using LLMs to solve planning problems. Several papers have demonstrated the potential of using LLMs for decision-making Sharma et al. (2022); Ren et al. (2023); Ichter et al. (2022); Singh et al. (2022). Some techniques that enable the use of LLMs as planners involve decomposing abstract tasks into concrete, atomic executable actions for an agent to perform Huang et al. (2022b, a); Sharma et al. (2022). Other approaches generate plan actions by scoring the possible next steps in a sequence of actions Ichter et al. (2022); Ren et al. (2023). When using LLMs to generate plans directly, the LLM is given a natural language representation of a planning problem with the goal of generating a plan. We refer to this line of work as “LLMs as planners.” One of the main findings is that LLMs have limited ability to generate and validate plans on their own, even for simple planning tasks Valmeekam et al. (2022); Silver et al. (2022); Valmeekam et al. (2023b).
Planner-Augmented LLMs.
Given LLMs’ poor performance in classical planning tasks Valmeekam et al. (2023a, 2022); Silver et al. (2022), new approaches extend LLMs by incorporating classical planning techniques Liu et al. (2023); Guan et al. (2023). Some techniques frame the problem as a machine translation task, transforming a natural language text description of a planning problem into PDDL to use with classical planners Liu et al. (2023); Guan et al. (2023); Xie et al. (2023). We refer to this line of work as “Planner-Augmented LLMs.” Other similar approaches translate natural language into alternative representations of the planning problem, such as finite state automata Yang et al. (2023a) or logic programming Yang et al. (2023b), to solve them.
Benchmarking LLMs on Planning Tasks.
Significant efforts have been made to develop benchmarks to assess the use of “LLMs as Planners.” PlanBench Valmeekam et al. (2023a) evaluates LLMs on various planning-related tasks, including plan generation, cost-optimal planning, plan verification, and others. Their work focuses on evaluating LLMs on their ability to generate plans. In contrast, Planetarium focuses on evaluating “Planner-Augmented LLMs”. Here, the goal is to determine whether an LLM can successfully translate natural language descriptions of planning problems into the correct PDDL representations of those problems. This is arguably a more natural use of LLMs, but it is harder to evaluate because the output is an internal representation and not the end result of the planning process.
Planning problem equivalence.
There is also work exploring techniques for determining the equivalence of planning problems, such as Shrinah et al. (2021). Chrpa et al. (2023) and Sievers et al. (2021) use a lifted graph representation of the planning domain, with the former specifically using graph isomorphism to check equivalence. Planetarium also employs graph isomorphism, but our focus is on evaluating the equivalence of PDDL problem instances rather than domains. Since domains and problem instances differ, the representation and isomorphism checks must also be adapted accordingly.
@@ 3.1 Classical Planning Problems
We use the set-theoretic form of classical planning problems Ghallab et al. (2004).
@@
  • $A$ is the set of actions. Each action $a$ is a triplet $({\texttt{pre}({a})},~{}{\texttt{add}({a})},~{}{\texttt{del}({a})}$, where ${\texttt{pre}({a})}$ is the set of preconditions that must be satisfied for $a$ to be executed, ${\texttt{add}({a})}$ is the set of positive effects that become true after executing $a$, and ${\texttt{del}({a})}$ is the set of negative effects that become false after executing $a$.
  • The transition function $\gamma(s,a)=(s-\texttt{del}(a))\cup\texttt{add}(a)$ models how the world changes by an action.
@@
A plan $\pi=<a_{1},a_{n},\ldots,a_{n}>$ is a sequence of actions that leads from $s_{i}$ to any state in $S_{g}$ following $\gamma$. This plan is a solution to $P$.
@@ 3.2 Planning Domain Definition Language
PDDL is a specialized language designed to provide a unified way to represent planning problems. It can represent various types of planning problems, including classical planning problems McDermott (2000); Gerevini and Long (2006). A PDDL planning problem consists of two files: the domain file, which describes the constant parts of the world model, and a problem file, which specifies particular planning scenarios using the world model outlined in the domain file.
The domain file specifies the set of possible actions $A$ and their preconditions and effects, which collectively define the transition function $\gamma$. The problem file defines the initial state $s_{i}$ and the goal propositions $g$. Finally, the set of proposition symbols $L$ is created by combining the predicates from the domain file with the objects defined in the problem file.
@@ 3.3 Scene Graphs
To compare different planning problems, we use scene graphs. A scene graph is a data structure commonly used in fields such as computer vision and graphics Johnson et al. (2015); Chang et al. (2023), rearrangement Ramachandruni et al. (2023), and planning to represent objects, their attributes, and the relationships among them. In our work, we define scene graphs as directed graphs with types and attributes for both nodes and edges. We represent a PDDL problem file with scene graphs as follows. We create one scene graph (the initial scene) for the initial state and another for the set of goal propositions (the goal scene). For every object, we create a node with an object type. Then, for every proposition that is listed in the problem file, we create a node with a proposition type. That node is given an attribute with the name of its predicate. Then, for each argument to the predicate in that proposition, we add an edge from the proposition node to the corresponding object. Each edge is given three attributes: the name of the predicate, the position of the argument (first, second, etc.), and whether it is defined in the initial state or the goal propositions. A scene graph is thus ($(\{O\cup P\},E)$, where $O$ is the set of nodes with object type, $P$ is the set of nodes with Proposition type, and $E$ is the set of edges.
We further define a problem graph as the combination of an initial scene and goal scene. Given $\textit{{SceneGraph}}_{\textit{{init}}}=(\{O\cup P_{\textit{{init}}}\},E_{\textit{{init}}})$ and $\textit{{SceneGraph}}_{\textit{{goal}}}=(\{O\cup P_{\textit{{goal}}}\},E_{\textit{{goal}}})$, then a problem graph merges the scene graphs such that $\textit{{ProblemGraph}}=(\{O\cup P_{\textit{{init}}}\cup P_{\textit{{goal}}}\},\{E_{\textit{{init}}}\cup E_{\textit{{goal}}}\})$. We define graph isomorphism on any of these graphs as an edge and type-preserving bijection between nodes, meaning that two graphs share a connectivity structure where all types and attributes match. See Appendix B for examples of diagrams of scene and problem graphs.
@@ 4 Evaluation Framework
In this section, we describe the design of our benchmark. Planetarium consists of two components: an algorithm that validates whether two PDDL problem files, a ground truth file and an LLM-generated file, are equivalent, and a curated dataset of planning tasks against which an LLM can be evaluated.
@@ 4.1 Planning Problem Equivalence
The first step to benchmarking PDDL generation is determining how to decide whether the generated code matches ground truth code. One might assume that checking if two PDDL problem files are equivalent is straightforward. However, the same goal state could be represented by many PDDL problem files, as shown in Figure 1.
Given these difficulties, we propose a definition of equivalence in terms of classical planning problems. The main idea is to find a bijective function between the sets of proposition symbols $L$ of the two problems such that makes the two problems equal. Our definition assumes that the transition function $\gamma$ is shared between the two problems. Our formal definition of equivalence between two planning problems follows.
@@
Two planning problems $P_{1}=(L^{1},S^{1},A^{1},\gamma,s_{i}^{1},g^{1})$ and $P_{2}=(L^{2},S^{2},A^{2},\gamma,s_{i}^{2},g^{2})$ with the same transition function $\gamma$ are equivalent if there exists a bijective function $f:L^{1}\to L^{2}$ such that:
@@
  4. 4.
    $\begin{array}[]{lcl}A^{2}=\{(&\{f(p):p\in\texttt{pre}(a)\}&,\\ &\{f(p):p\in\texttt{add}(a)\}&,\\ &\{f(p):p\in\texttt{del}(a)\}&):a\in A^{1}\}.\end{array}$
Figure 2: A demonstration of how fullySpecify fills in gaps of the goal state of a planning problem.
That is, all predicates across the states, action definitions, initial state, goal specification, and actions of one problem are replaced with their corresponding predicates from the other.
This definition is not directly usable for checking equivalence between two PDDL problem files because it relies on finding a bijection between the sets $L$ of each problem, and PDDL does not define these sets directly. To build the set $L$ with PDDL, one needs to use the predicates in the domain file and the objects in the problem file, instantiating each predicate with all possible objects they might take. We assume the entire PDDL domain is shared between the problems. This fact entails that the predicates will be shared, making it necessary to look for bijective functions for PDDL only over the objects. The challenge is that each PDDL problem file can correspond to many pairs of initial and goal states because the set of goal propositions makes an open-world assumption. That set can leave implicit trivial propositions that are necessarily true and therefore do not change the underlying planning problem. We must therefore fully specify the goal, meaning that we identify all propositions that are true in all reachable goal states when starting from the initial state.
Our algorithm for checking equivalence is summarized in Algorithm 1. First, we transform each set of initial state and goal state propositions into scene graphs. Second, we fully specify the goal scene graphs by adding all trivially true edges. Finally, we join the initial state scene graph with each goal state graph to create problem graphs and look for a bijection between objects such that the problem graphs are isomorphic.
@@
1:function equivalent($\text{P}_{a},\text{P}_{b},\text{isPlaceholder}$)
@@
4:     if $\text{canDoFast}(\text{s}_{i,a},\text{g}_{a},\text{s}_{i,b},\text{g}_{b})$ then
5:         return $\text{fastEquivalent}(\text{s}_{i,b},\text{g}_{a},\text{s}_{i,a},\text{g}_{b})$
6:     end if
7:     $\text{g}^{\star}_{a}\leftarrow\text{fullySpecify}(\text{s}_{i,a},\text{g}_{a})$
8:     $\text{g}^{\star}_{b}\leftarrow\text{fullySpecify}(\text{s}_{i,b},\text{g}_{b})$
9:     if isPlaceholder then
@@
[…]
13:     else
14:         $\text{p}_{a}\leftarrow\text{join(}\text{s}_{i,a},\text{g}^{\star}_{a}\text{)}$
15:         $\text{p}_{b}\leftarrow\text{join(}\text{s}_{i,b},\text{g}^{\star}_{b}\text{)}$
16:         return $\text{isIsomorphic}(\text{p}_{a},\text{p}_{b})$
17:     end if
18:end function
Transform to Scene Graphs.
@@
From each PDDL problem file, we generate two scene graphs: one for the initial state and another for the goal state (lines 2–3). For each transformation, we first create an object node for each object. Then, for each proposition in the collection, we create a new proposition node and create edges between the proposition node and the objects it takes as arguments. Edge attributes denote argument order, predicate type, and whether the proposition is in an initial or goal scene. See Section 3.3 for details.
Check Easy Cases.
@@
For speed, we check several cases and return early if we can (lines 4–6). First, if the number of object nodes in each graph is not equal, the problems cannot be equivalent. Second, if the initial scenes are not isomorphic, then the problems cannot be equivalent. Finally, if the problem graphs, composed of the initial and goal scenes, are isomorphic, then the problems are equivalent.
Table 1: A breakdown of the Planetarium dataset by the level of abstractness of the text description, number of propositions in the ground truth PDDL problem (size), and train/test split. Each instance consists of a ground truth PDDL problem description paired with a text description.
Domain Total Abstractness Num. of Ground Truth Propositions (Size) Split
Explicit to Abstract to
Explicit Abstract Explicit Abstract 1-20 21-40 41-60 61-80 >80 Train Test
Blocks World 85,605 22,595 20,152 20,152 22,706 963 10,011 47,153 25,766 1,712 80,612 4,993
Gripper 54,141 13,683 10,565 12,466 17,427 2,334 13,640 16,405 20,325 1,437 49,363 4,778
Floor Tile 6,172 2,289 797 793 2,293 706 2,114 2,224 404 724 0 6,172
Fully Specify the Goal Scenes.
@@
If the input is not an easy case, then we have to reason about the sets of goal states defined by the goal scenes (Figure 2). Condition 3 of Definition 2 requires that the sets of goal states that are consistent with the given goal propositions be equal after substituting matching propositions. Since PDDL uses the open-world assumption for goals, we have to identify all propositions that are true in all reachable goal states when starting from the initial state for each problem. The function fullySpecify finds all such propositions and adds them to the goal scenes as additional proposition nodes with the corresponding edges and attributes (lines 7–8). We show how to implement fullySpecify efficiently for the three domains in Planetarium in Appendix C.
PDDL without Object Identity.
@@
Our algorithm runs in two modes, depending on the type of problem it is checking. Sometimes we want to compare a generated PDDL problem file with many ground truth problem files and see if the generated file matches any one of them. For example, if the natural language description says “make a tower of height 3,” the specific blocks to use are unspecified, and any permutation of blocks that builds a tower of height 3 should be considered correct. Concretely, we want to treat the objects in the PDDL goal states as placeholders and accept any permutation of them. We check this condition when isPlaceholder is True (determined by the problem type) simply by checking isomorphism between initial and goal scenes separately rather than combining them into problem graphs (lines 10–12).
PDDL with Object Identity.
@@
If isPlaceholder is False, we want to check whether the problem files are precisely equivalent under Definition 2, meaning that the objects in the initial scenes correspond to the same objects in the goal scenes as well. We first join the corresponding initial and scene graphs into problem graphs as described in Section 3.3 (lines 14–15). Then we check whether the two problem graphs are isomorphic (line 16).
@@
Equivalent($P_{a}$, $P_{b}$, False) returns True if and only if the PDDL problem files $P_{a}$ and $P_{b}$ represent equivalent planning problems under Definition 2. Equivalent($P_{a}$, $P_{b}$, True) returns True if and only if $P_{a}$ represents a planning problem that is equivalent to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state.